OpenStudy (anonymous):

If you put $2,000 in a savings account that pays 6% interest compounded continuously, how much money will you have in your account in 4 years? Assume you make no additional deposits or withdrawals.

OpenStudy (anonymous):

First, you should find out whats the 6% of 2000, easy way: 2000/100*6=?. Then you should multiply by the number of years 4(2000/100*6)=? Then you should add up these numbers 2000+4(2000/100*6)=?

OpenStudy (anonymous):

but i think really you could just use A=pe^rt which is A=(2000)e^(.06*4)

OpenStudy (anonymous):

=2542.498

OpenStudy (anonymous):

2000+4(2000/100*6)=2000+4*120=2480